marc goes to the store with exactly $1.00 in change. He has at least one of each coin less than a half-dollar coin, but does not have a half dollar coin. What is the most amount of coins he could have? What is the least amount of coins could he have?

Answer by htmentor(1343)

marc goes to the store with exactly $1.00 in change. He has at least one of each coin less than a half-dollar coin, but does not have a half dollar coin. What is the most amount of coins he could have? What is the least amount of coins could he have?
=====================================
He has at least one penny, nickel, dime and quarter. If he has one quarter, one dime, one nickel and all the rest pennies,
this would give him the most amount of coins.
25 + 10 + 5 + p = 100
where p = the number of pennies
This gives p = 60
So the most amount of coins = 63
For the least amount, we start with 3 quarters. That leaves 25 cents. Since we need at least one nickel and one penny,
we can only use one dime. This leaves 15 cents. So we need two nickels and five pennies.
So the least amount of coins = 3 + 1 + 2 + 5 = 11